How would I add the following MS curves together to achieve the following MS(t) curve, as shown in the graph?

MS(a)=200-10e
MS(b)=80-e

To give a MS(t) curve of :
200-10e when e<12
100-(10/11)e when e>=12

200 MST MSB 20 80 100 E
Answer #1

When MS is between 200 to 80, MSb=0,

When MSa =80 , 80=200-10e, e= 120/10=12

So for e<12, 200-10e

as only MSa is possible

For e≥12, We need horizontal summation of MSa and MSb.

So write the equation in terms of e and then add them.

MSa=200-10e

ea= 20-0.1MSa

MSb=80-e

eb=80-MSb

ADD ea and eb.

ea+eb=20-0.1MS+80-MS= 100-1.1MS=e

MS=(100-e)/1.1

So, for e≥12 ,MS=(100-e)/1.1

If it helps kindly upvote

For doubts comment below